i need help finding the peak area
Aeq. Operator : SYSTEM Aca. Instrument : 6890 GC Injection Date Seq. Line Location Vial 9 Inj Inj Volume 0.2 ul C:\Chem3212\DATA\ RH3-RH2-BS1-819\FRAC- DIST 2019-06-07 15-51-391FRAC-D1ST 12 6/7/2019 5:58:29 PM Sequence File .S Method C:\CHEM32\2\DATAARH3-RH2-HS1-S191FRAC-DIST 2019-06-07 15-51-39\FRACDIST. M (Sequence Method) 6/7/2019 3:51:39 PM by SYSTEM Fractional Distillation of Ethyl acetate and Butyl acetate Last changed Method Info FID1 A, (FRACDISTO00012 D) pA buly acetuic einyl acciaic 80000 60000 40000 20000 0 mini 2.5 1.5 0.5 Area Percent Report Signal 1.0000 Sorted By Multiplier Dilution Use Multiplier & Dilution Factor with ISTDS 1.0000
nTST 2019-06-07 15 aFile Cr\CHEM3212DATAARH3-R12-8S1-819VFRAC-DIST 2019-06-07 15-51-39ArRACDISTOOO011.D Eple Name: Acq. Operator : SYSTEM Seg. Line t Location Vial 8 Inj i Ind Volume: 0.2 pl 11 Acq. Instrument : 6890 GC Injection Date : 6/7/2019 5:47:02 PM Sequence File : C:\Chen32\2\DATA\RH3-R82-HS1-819 FRAC -DIST 2019-06-07 15-51-39\ FRAC - DIST Method : CINCHEM3212\ DATA\ RM3-R82-HS1-s19\FRAC -DIST 2019-06-07 15-51-391FRACDIST M (Sequence Method) : 6/7/2019 3:51:39 PM by SYSTEM Fractional Distillation of Ethyl acetate and Butyl acetate Last changed Method Info FID1 A, (FRACDSTOO0011 0) cinyl ucctulc 175000- 150000- 125000- 100000 75000- bulyi ucelate 50000- 25000- 0- min 15 25 0.5 Area Percent Report Signal 1.0000 1.0000 Sorted By Multiplier Dilution Use Multip1ier & Dilution Factor with ISTDS
1 Fractional Distillation of Ethyl acetate d aMue ct e FID1 A (FRACOISTOO0010.D)y elhyl acelate 200000 175000 150000 125000- 100000- 75000- 50000 bulyl acclule 25000 0 miri 1.5 25 0.5 Area Percent Report Signal 1.0000 1.0000 Sorted By Multiplier Dilution Use Multiplier & Dilution Factor with ISTDs
Answer #1

1. The peak area of ethyl acetate = {87500/(87500+97500)}*100 = 47.3%

Now, the peak area of butyl acetate = 100 - 47.3 = 52.7%

2. The peak area of ethyl acetate = {190000/(42500+190000)}*100 = 81.7%

Now, the peak area of butyl acetate = 100 - 81.7 = 18.3%

3. The peak area of ethyl acetate = {205000/(17500+205000)}*100 = 92.1%

Now, the peak area of butyl acetate = 100 - 92.1 = 7.9%